Eight fair coins were flipped. 21. How many possible outcomes there will be, if the order of coins are considered? 22. How many possible outcomes with exactly 3 heads? 23. What is the probability of getting a result with exactly 3 heads? (Round to 4 decimal places, if needed.) What is the probability of getting a result with less than 2 head? (Round to 4 decimal places, if needed.) 24. 25. P(A)-0.35, P(B)-0.58, and P(A and B)-0.26, What is P(A or B)? (Round to 2 decimal places, if needed.) 26. P(A) 0.75, P(B)-0.36, and PCA and B)-0.25, What is P(A or B)? (Round to 2 decimal places, if needed.)
Answer #1

Solution:-

21. There are 2^8 = 256 possible outcomes.

22. C(8,3) = 56 possible outcomes with exactly 3 heads.

23. by using data : n = 8, X = 3 , p = 0.5

P(X = 3) = 8C3*(0.5^3)*(0.5^5) = 0.2187

24. P(X < 2) = 0.0352

25. P(A or B) = P(A) + P(B) - P(A and B)
= 0.35 + 0.58 - 0.26
= 0.67

26. P(A or B) = 0.75 + 0.36 - 0.25 = 0.86